# On the Design of Expressive and Trainable Pulse-based Quantum Machine Learning Models

### v1

@@ Abstract
Pulse-based Quantum Machine Learning (QML) has emerged as a novel paradigm in quantum artificial intelligence due to its exceptional hardware efficiency. For practical applications, pulse-based models must be both expressive and trainable. Previous studies suggest that pulse-based models under dynamic symmetry can be effectively trained, thanks to a favorable loss landscape that has no barren plateaus. However, the resulting uncontrollability may compromise expressivity when the model is inadequately designed. This paper investigates the requirements for pulse-based QML models to be expressive while preserving trainability. We present a necessary condition pertaining to the system’s initial state, the measurement observable, and the underlying dynamical symmetry Lie algebra, supported by numerical simulations. Our findings establish a framework for designing practical pulse-based QML models that balance expressivity and trainability.
1 INTRODUCTION
@@
Quantum computing holds substantial promise for significantly improving the efficiency of both training and inference in machine learning tasks [1, 2, 3, 4]. However, in the noisy intermediate-scale quantum (NISQ) era, the lack of error correction necessitates that quantum machine learning (QML) collaborates with classical computers [5, 6] through variational quantum models. In this framework, quantum models execute the inference task, while the classical computer manages the training.
@@
In the literature, variational QML models predominantly employ gate-based parameterized quantum circuits, also known as quantum neural networks (QNNs) [7, 8]. Since all gates are ultimately implemented through control pulses on the relevant physical qubits, recent proposals suggest that QML models may be parameterized directly by these continuous-time control pulses instead of discrete-time gate parameters [9, 10, 11, 12, 13, 14, 15, 16]. In contrast to gate-based models, pulse-based models are grounded in physical systems without abstraction, making them more accessible to engineers due to their straightforwardness and hardware efficiency [17, 18, 19, 20, 21].
@@
In principle, within the finite coherence time, pulse-based QML models are more expressive compared to their gate-based counterparts because they function as "infinitely" deep data-re-uploading QNNs [22]. This advantage has been demonstrated in our previous study [22], where we also proved that pulse-based models achieve full expressivity (i.e., the ability to approximate arbitrary functions [23, 24, 25, 26]) when the underlying quantum system is ensemble controllable. This criterion provides a control-oriented framework for designing pulse-based models, analogous to the design of universal quantum computers [27]. Nonetheless, in systems comprising a large number of qubits, the requirement for controllability becomes excessively stringent as it inevitably leads to barren plateaus within the loss landscape [28, 29]. This phenomenon results in an exponential decay of loss gradients (or, equivalently, the loss variance [30, 31]) in the asymptotic temporal limit [32] and the large qubit-number limit, thus rendering optimization within large-scale systems impracticable [8, 33, 34, 35, 36, 37, 38]. A similar trade-off between expressivity and trainability is observed in gate-based models, where the loss function predominantly concentrates as the circuit attains greater expressivity [33, 39, 40].
@@
To alleviate the conflict between expressivity and trainability, controllability must be sacrificed in the design of pulse-based models. Larocca et al. [28] reveals that trainability is contingent on the degree of controllability characterized by the size and symmetry of its dynamical Lie algebra. The findings indicate that a symmetry-restricted model has the potential to circumvent barren plateaus, thus providing a practical diagnostic approach for the design of QML models. Building upon this insight, a Lie-algebraic theory for barren plateaus has been developed, based on which an exact variance formula is derived for the loss function. It is demonstrated that the variance scales inversely with the dimension of the dynamical Lie algebra, provided that either the initial state or observable is contained within the Lie algebra [41, 42]. Diaz et al. [43] extends beyond the Lie-algebraic framework by analyzing parameterized matchgate circuits. By decomposing the operator space into Lie group modules, they derive an exact variance formula for loss functions applicable to arbitrary initial states and observables, demonstrating that barren plateaus can emerge from generalized global operators and are dependent on module dimensions rather than solely on the dimension of the dynamical Lie algebra.
@@
The existing literature provides a Lie algebraic framework for evaluating the trainability of uncontrollable pulse-based QML models induced by dynamic symmetry, but the expressivity of these models remains ambiguous. In the present study, we demonstrate that an ill-suited selection of either the initial state or the measurement observable can result in the loss of expressivity, and conduct a systematic analysis on the expressivity of pulse-based QML models within the context of dynamic symmetry. In contrast to the Fourier-series analysis typically used to assess the expressivity of data-reuploading gate-based models [44, 45, 46, 47, 48], we utilize the polynomial expansion of nonlinear functions epitomized by the pulse-based model based on Dyson-series analysis (also known as Fliess-series in control theory for input-output analysis of nonlinear systems [49]). This analysis yields a Lie algebraic tool for the design of expressive pulse-based models with dynamic symmetry.
@@
The structure of the paper is as follows: Section 2 introduces essential background on pulse-based QML models and the existing results on the trainability under dynamic symmetry. Section 3 analyzes the expressivity of pulse-based models through the Dyson-series expansion from which a necessary condition is derived. In Section 4, numerical experiments are conducted to demonstrate the expressivity of pulse-based models with dynamic symmetry, and the balance between expressivity and trainability. Finally, Section 5 draws the conclusion and future perspectives.
2 PULSE-BASED QML MODELS with dynamical symmetry
@@
Figure 1(b) illustrates that the physical implementation of the quantum circuit model consistently compiles individual quantum gates into time-varying control pulses on the physical qubits. Let $x_{k}(t)$ and $\theta_{k}(t)$ represent the corresponding data-encoding and training control pulses applied to the $k$th qubit. The QML process can be adjusted by $\Theta=\{\theta_{1}(t),\cdots,\theta_{n}(t),\leavevmode\nobreak\ t\in[0,T]\}$, resulting in a pulse-based model implemented at the physical level. Furthermore, as shown in Fig. 1(c), the pulses can be continuously and simultaneously applied without waiting for one another, allowing for compression to reduce inference time. Such generalized pulse-based model is free from a prior circuit design, featuring more tunable parameters within the finite coherence time, and can achieve greater expressive power than gate-based models [22].
@@
in which the encoding Hamiltonian
@@
is introduced to enhance the expressivity of the pulse-based model. The time-variant functions $\theta_{1}(t),\cdots,\theta_{m+p}(t)$ are control pulses that can be trained to fit the dataset. Similar to gate-based models, the output is inferred through the expectation-value measurement
@@
with $\Theta=\{\theta_{1}(t),\cdots,\theta_{m+p}(t),\leavevmode\nobreak\ t\in[0,T]\}$. In practice, certain control pulses may be constant or exhibit limited tunability due to hardware constraints.
@@
be its decomposition as a direct sum of commuting ideals, where $\mathfrak{c}$ is the center of $\mathfrak{g}$ and $\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{k}$ is the semi-simple part. Assume that either the initial state $|\rho=\ket{\psi_{0}}\bra{\psi_{0}}$ or the observable $M$ belongs to the Lie algebra $\mathfrak{g}$ (a relaxed condition can be found in [43]), and the QNN is sufficiently deep (i.e., a sufficiently long time duration $T$ for pulse-based QNN) to establish a $2$-design over the Lie group $e^{\mathfrak{g}}$ [41, 42]). The variance of $f(\textbf{x},\Theta)$ can be estimated as follows [41, 29]
@@
denotes the sqaure norm of the orthogonal projection [with respect to the Hilbert-Schmidt inner product $\langle A,B\rangle\overset{\rm def}{=}{\rm Tr}(A^{\dagger}B)$] of $H$ onto $\mathfrak{g}_{\mathbb{C}}$ (the complexification of $\mathfrak{g}$). $\{B_{j}\}^{\rm dim(\mathfrak{g})}_{j=1}$ is an orthonormal basis for $\mathfrak{g}$.
The formula (6) indicates that the model becomes untrainable when the underlying control system is fully controllable, because in such case ${\rm dim}(\mathfrak{g})=4^{n}$ with $\mathfrak{g}=\mathfrak{su}(2^{n})$, yielding exponentially decreasing variance as $n$ increases. Therefore, non-trivial dynamic symmetry must be introduced to constrain the quantum state evolution to a manifold whose dimension does not increase exponentially with $n$.
3 EXPRESSIVITY OF QML MODELS in presence of dynamical symmetry
The pulse-based model (2) is said to be expressive if, for any function $f_{0}(\textbf{x})$ and a specified error threshold $\epsilon>0$, there exists a suitable control pulse $\Theta$ and a scaling factor $\theta_{0}$ such that the measurement-induced function (5) can approximate the function with the desired precision, i.e., $\|\theta_{0}f(\textbf{x},\Theta)-f_{0}(\textbf{x})\|<\epsilon$.
@@ 3 EXPRESSIVITY OF QML MODELS in presence of dynamical symmetry
In the following, we demonstrate that such uncontrollable pulse-based models can be expressive when the initial state and measurement observable are appropriately selected. To derive the required condition, we expand the output function (5) into a polynomial series of $\mathbf{x}$. For convenience, we first rewrite (2) as a Liouville equation:
@@
We then apply the Dyson series expansion (also known as the Fliess expansion in control theory literature [49]):
@@
By substituting (8) and expanding the series into polynomial terms of $x_{1},\cdots,x_{m}$, we obtain the following polynomial expansion:
@@
The polynomial expansion clearly demonstrates the inherent nonlinearity in the model induced by a continuous-limit data-reuploading mechanism [22]. It is evident that the pulse-based model can approximate any continuous function if the coefficients $\{C_{k_{1},\cdots,k_{m}}(\Theta)\}$ in the polynomial series are fully adjustable by tuning the control pulses $\theta_{1}(t),\cdots,\theta_{m+p}(t)$. According to (9), the adjustability of these coefficients depends on both the control pulses, through $c_{j_{1}\cdots j_{n}}(\Theta)$ terms, and the initial state and observable through $\langle\psi_{0}|\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M|\psi_{0}\rangle$ terms.
Ideally, if the coefficients $c_{j_{1}\cdots j_{n}}(\Theta)$ are mutually independent and fully adjustable, the model is expressive provided that, for each $(k_{1},\cdots,k_{m})$, there exist some $(j_{1},\cdots,j_{n})\in\mathbb{J}_{k_{1},\cdots,k_{m}}$ such that $\langle\psi_{0}|\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M|\psi_{0}\rangle\neq 0$. Unfortunately, the mutual independence of $c_{j_{1}\cdots j_{n}}(\Theta)$ does not hold, as it can be negated by the counterexample
$c_{j_{1}j_{2}}(\Theta)+c_{j_{2}j_{1}}(\Theta)=c_{j_{1}}(\Theta)c_{j_{2}}(\Theta),\quad\forall\leavevmode\nobreak\ i,j.$
@@
To facilitate the validation of this condition, we denote by
$\mathcal{S}_{k_{1},\cdots,k_{m}}=\{\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M,\leavevmode\nobreak\ (j_{1},\cdots,j_{n})\in\mathbb{J}_{k_{1},\cdots,k_{m}}\}$
@@
$\begin{split}\mathfrak{M}(\mathcal{S})=&\leavevmode\nobreak\ {\rm span}\{\mathcal{L}_{\alpha_{1}}\cdots\mathcal{L}_{\alpha_{\ell}}X,\leavevmode\nobreak\ X\in\mathcal{S},\\ &\quad m+1\leq\alpha_{i}\leq m+p,\leavevmode\nobreak\ \ell\in\mathbb{N}\}\end{split}$ (11)
the submodule generated by the operator set $\mathcal{S}$ with the control Liouvillians. By definition, it is apparent that
@@
because $\mathbb{J}_{0,\cdots,0}$ contains only tuples $(j_{1}\cdots,j_{n})$ with $m+1\leq j_{k}\leq m+p$. Starting from $\mathcal{S}_{0,\cdots,0}$, we can recursively construct $\mathcal{S}_{k_{1},\cdots,k_{m}}$ as follows:
@@
Theorem 1 provides a method for identifying candidate pulse-based QML models that are both trainable and expressive. Although the sufficiency for expressivity cannot be rigorously proven, subsequent numerical simulations will demonstrate that models meeting this condition tend to be highly expressive. This is attributed to the infinite number of $c_{j_{1}\cdots j_{n}}(\Theta)$ terms in (9), which offer considerable flexibility even if they are not mutually independent.
4 Numerical Simulations
In this section, we demonstrate through numerical simulations the expressivity and trainability of pulse-based QML models under selected dynamic symmetries.
Given a training dataset $\{(\textbf{x}^{(k)},y^{(k)})\}$ that contain $N$ samples, the model is trained by minimizing the loss function

### v2

@@ Abstract
Pulse-based Quantum Machine Learning (QML) has emerged as a novel paradigm in quantum artificial intelligence due to its exceptional hardware efficiency. For practical applications, pulse-based models must be both expressive and trainable. Previous studies suggest that pulse-based models under dynamic symmetry can be effectively trained, thanks to a favorable loss landscape that avoids barren plateaus. However, the resulting uncontrollability may compromise expressivity when the model is inadequately designed. This paper investigates the requirements for pulse-based QML models to be expressive while preserving trainability. We establish a necessary condition pertaining to the system’s initial state, the measurement observable, and the underlying dynamical symmetry Lie algebra, supported by numerical simulations. Our findings provide a framework for designing practical pulse-based QML models that balance expressivity and trainability.
I INTRODUCTION
@@
Quantum computing holds substantial promise for improving the efficiency of both training and inference in machine learning tasks [1, 2, 3, 4]. However, in the noisy intermediate-scale quantum (NISQ) era, the lack of error correction necessitates that quantum machine learning (QML) collaborates with classical computers [5, 6] through variational quantum models. In this framework, quantum models execute the inference task, while the classical computer manages the training.
@@
Variational QML models predominantly employ gate-based parameterized quantum circuits, also known as quantum neural networks (QNNs) [7, 8]. Since all gates are ultimately implemented through control pulses on the relevant physical qubits, recent proposals suggest that QML models may be parameterized directly by these continuous-time control pulses instead of discrete-time gate parameters [9, 10, 11, 12, 13, 14, 15, 16]. In contrast to gate-based models, pulse-based models are grounded in physical systems without abstraction, making them more accessible to engineers due to their straightforwardness and hardware efficiency [17, 18, 19, 20, 21].
@@
Within the finite coherence time, pulse-based QML models are more expressive compared to their gate-based counterparts because they function as "infinitely" deep data-re-uploading QNNs [22]. We demonstrated this advantage in our previous study [22], where we also proved that pulse-based models achieve full expressivity (i.e., the ability to approximate arbitrary functions [23, 24, 25, 26]) when the underlying quantum system is ensemble controllable. This criterion provides a control-oriented framework for designing pulse-based models, analogous to the design of universal quantum computers [27]. Nonetheless, in systems comprising a large number of qubits, the requirement for controllability becomes excessively stringent as it inevitably leads to barren plateaus within the loss landscape [28, 29]. This phenomenon results in an exponential decay of loss gradients (or, equivalently, the loss variance [30, 31]) in the asymptotic temporal limit [32] and the large qubit-number limit, thus rendering optimization within large-scale systems impracticable [8, 33, 34, 35, 36, 37, 38]. A similar trade-off between expressivity and trainability is observed in gate-based models, where the loss function predominantly concentrates as the circuit attains greater expressivity [33, 39, 40].
@@
To alleviate the conflict between expressivity and trainability, controllability must be sacrificed in the design of pulse-based models. Larocca et al. [28] revealed that trainability is contingent on the degree of controllability characterized by the size and symmetry of its dynamical Lie algebra. Their findings indicate that a symmetry-restricted model can potentially circumvent barren plateaus, thus providing a practical diagnostic approach for QML model design. Building upon this insight, a Lie-algebraic theory for barren plateaus has been developed, based on which an exact variance formula is derived for the loss function. It is demonstrated that the variance scales inversely with the dimension of the dynamical Lie algebra, provided that either the initial state or observable is contained within the Lie algebra [41, 42]. Diaz et al. [43] extended beyond the Lie-algebraic framework by analyzing parameterized matchgate circuits. By decomposing the operator space into Lie group modules, they derived an exact variance formula for loss functions applicable to arbitrary initial states and observables, demonstrating that barren plateaus can emerge from generalized global operators and depend on module dimensions rather than solely on the dimension of the dynamical Lie algebra.
@@
The existing literature provides a Lie algebraic framework for evaluating the trainability of uncontrollable pulse-based QML models induced by dynamic symmetry, but the expressivity of these models remains ambiguous. In this study, we demonstrate that an ill-suited selection of either the initial state or the measurement observable can result in the loss of expressivity, and conduct a systematic analysis of the expressivity of pulse-based QML models within the context of dynamic symmetry. In contrast to the Fourier-series analysis typically used to assess the expressivity of data-reuploading gate-based models [44, 45, 46, 47, 48], we utilize the polynomial expansion of nonlinear functions epitomized by the pulse-based model based on Dyson-series analysis (also known as Fliess-series in control theory for input-output analysis of nonlinear systems [49]). This analysis yields a Lie algebraic tool for designing expressive pulse-based models with dynamic symmetry.
The main contributions of this work are: (i) establishing a necessary condition for expressivity in pulse-based QML models under dynamic symmetry; (ii) demonstrating through numerical simulations that models satisfying this condition can achieve a balance between expressivity and trainability; and (iii) providing a framework for designing practical pulse-based QML models.
@@
The structure of the paper is as follows: Section II introduces essential background on pulse-based QML models and existing results on trainability under dynamic symmetry. Section III analyzes the expressivity of pulse-based models through Dyson-series expansion and derives a necessary condition. In Section IV, numerical experiments demonstrate the expressivity of pulse-based models with dynamic symmetry and the balance between expressivity and trainability. Finally, Section V presents conclusions and future perspectives.
II PULSE-BASED QML MODELS WITH DYNAMICAL SYMMETRY
@@
Figure 1(b) illustrates that the physical implementation of the quantum circuit model compiles individual quantum gates into time-varying control pulses on the physical qubits. Let $x_{k}(t)$ and $\theta_{k}(t)$ represent the corresponding data-encoding and training control pulses applied to the $k$th qubit. The QML process can be adjusted by $\Theta=\{\theta_{1}(t),\cdots,\theta_{n}(t),\penalty 10000\ t\in[0,T]\}$, resulting in a pulse-based model implemented at the physical level. Furthermore, as shown in Fig. 1(c), the pulses can be continuously and simultaneously applied without waiting for one another, allowing for compression to reduce inference time. This generalized pulse-based model is free from prior circuit design, features more tunable parameters within the finite coherence time, and can achieve greater expressive power than gate-based models [22].
@@
where the encoding Hamiltonian
@@
enhances the expressivity of the pulse-based model. The time-variant functions $\theta_{1}(t),\cdots,\theta_{m+p}(t)$ are control pulses that can be trained to fit the dataset. Similar to gate-based models, the output is inferred through the expectation-value measurement
@@
with $\Theta=\{\theta_{1}(t),\cdots,\theta_{m+p}(t),\penalty 10000\ t\in[0,T]\}$. In practice, certain control pulses may be constant or exhibit limited tunability due to hardware constraints.
@@
be its decomposition as a direct sum of commuting ideals, where $\mathfrak{c}$ is the center of $\mathfrak{g}$ and $\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{k}$ is the semi-simple part. Assuming that either the initial state $|\rho=\ket{\psi_{0}}\bra{\psi_{0}}$ or the observable $M$ belongs to the Lie algebra $\mathfrak{g}$ (a relaxed condition can be found in [43]), and the QNN is sufficiently deep (i.e., a sufficiently long time duration $T$ for pulse-based QNN) to establish a $2$-design over the Lie group $e^{\mathfrak{g}}$ [41, 42]), the variance of $f(\textbf{x},\Theta)$ can be estimated as follows [41, 29]
@@
denotes the square norm of the orthogonal projection [with respect to the Hilbert-Schmidt inner product $\langle A,B\rangle\overset{\rm def}{=}{\rm Tr}(A^{\dagger}B)$] of $H$ onto $\mathfrak{g}_{\mathbb{C}}$ (the complexification of $\mathfrak{g}$). $\{B_{j}\}^{\rm dim(\mathfrak{g})}_{j=1}$ is an orthonormal basis for $\mathfrak{g}$.
Equation (6) indicates that the model becomes untrainable when the underlying control system is fully controllable, because in such cases ${\rm dim}(\mathfrak{g})=4^{n}$ with $\mathfrak{g}=\mathfrak{su}(2^{n})$, yielding exponentially decreasing variance as $n$ increases. Therefore, non-trivial dynamic symmetry must be introduced to constrain the quantum state evolution to a manifold whose dimension does not increase exponentially with $n$.
III EXPRESSIVITY OF QML MODELS IN PRESENCE OF DYNAMICAL SYMMETRY
The pulse-based model (2) is expressive if, for any function $f_{0}(\textbf{x})$ and a specified error threshold $\epsilon>0$, there exists a suitable control pulse $\Theta$ and a scaling factor $\theta_{0}$ such that the measurement-induced function (5) approximates the function with the desired precision, i.e., $\|\theta_{0}f(\textbf{x},\Theta)-f_{0}(\textbf{x})\|<\epsilon$.
@@ III EXPRESSIVITY OF QML MODELS IN PRESENCE OF DYNAMICAL SYMMETRY
We now demonstrate that uncontrollable pulse-based models can be expressive when the initial state and measurement observable are appropriately selected. To derive the required condition, we expand the output function (5) into a polynomial series of $\mathbf{x}$. For convenience, we first rewrite (2) as a Liouville equation:
@@
We apply the Dyson series expansion (also known as the Fliess expansion in control theory literature [49]):
@@
Substituting (8) and expanding the series into polynomial terms of $x_{1},\cdots,x_{m}$, we obtain the polynomial expansion:
@@
This polynomial expansion demonstrates the inherent nonlinearity in the model induced by a continuous-limit data-reuploading mechanism [22]. The pulse-based model can approximate any continuous function if the coefficients $\{C_{k_{1},\cdots,k_{m}}(\Theta)\}$ in the polynomial series are fully adjustable by tuning the control pulses $\theta_{1}(t),\cdots,\theta_{m+p}(t)$. According to (9), the adjustability of these coefficients depends on both the control pulses, through $c_{j_{1}\cdots j_{n}}(\Theta)$ terms, and the initial state and observable through $\langle\psi_{0}|\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M|\psi_{0}\rangle$ terms.
If the coefficients $c_{j_{1}\cdots j_{n}}(\Theta)$ were mutually independent and fully adjustable, the model would be expressive provided that, for each $(k_{1},\cdots,k_{m})$, there exist some $(j_{1},\cdots,j_{n})\in\mathbb{J}_{k_{1},\cdots,k_{m}}$ such that $\langle\psi_{0}|\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M|\psi_{0}\rangle\neq 0$. Unfortunately, the mutual independence of $c_{j_{1}\cdots j_{n}}(\Theta)$ does not hold, as demonstrated by the counterexample
$c_{j_{1}j_{2}}(\Theta)+c_{j_{2}j_{1}}(\Theta)=c_{j_{1}}(\Theta)c_{j_{2}}(\Theta),\quad\forall\penalty 10000\ i,j.$
Nevertheless, we can derive a necessary condition from this analysis. Intuitively, if for a particular monomial $x_{1}^{k_{1}}\cdots x_{m}^{k_{m}}$, all the coefficients $\langle\psi_{0}|\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M|\psi_{0}\rangle$ vanish, then the model cannot represent that monomial regardless of the control pulses chosen. This implies that the model loses expressivity for that monomial. Therefore, to approximate arbitrary functions, the model must be able to represent all monomials, which requires that for every monomial, at least one of the coefficients $\langle\psi_{0}|\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M|\psi_{0}\rangle$ is non-zero.
@@
To facilitate validation of this condition, we denote by
$\mathcal{S}_{k_{1},\cdots,k_{m}}=\{\mathcal{L}_{j_{1}}\cdots\mathcal{L}_{j_{n}}M,\penalty 10000\ (j_{1},\cdots,j_{n})\in\mathbb{J}_{k_{1},\cdots,k_{m}}\}$
@@
$\begin{split}\mathfrak{M}(\mathcal{S})=&\penalty 10000\ {\rm span}\{\mathcal{L}_{\alpha_{1}}\cdots\mathcal{L}_{\alpha_{\ell}}X,\penalty 10000\ X\in\mathcal{S},\\ &\quad m+1\leq\alpha_{i}\leq m+p,\penalty 10000\ \ell\in\mathbb{N}\}\end{split}$ (11)
the submodule generated by the operator set $\mathcal{S}$ with the control Liouvillians. By definition,
@@
because $\mathbb{J}_{0,\cdots,0}$ contains only tuples $(j_{1}\cdots,j_{n})$ with $m+1\leq j_{k}\leq m+p$. Starting from $\mathcal{S}_{0,\cdots,0}$, we recursively construct $\mathcal{S}_{k_{1},\cdots,k_{m}}$ as follows:
@@
Theorem 1 provides a method for identifying candidate pulse-based QML models that are both trainable and expressive. Although sufficiency for expressivity cannot be rigorously proven, subsequent numerical simulations demonstrate that models meeting this condition tend to be highly expressive. This is attributed to the infinite number of $c_{j_{1}\cdots j_{n}}(\Theta)$ terms in (9), which offer considerable flexibility even if they are not mutually independent.
IV NUMERICAL SIMULATIONS
We demonstrate through numerical simulations the expressivity and trainability of pulse-based QML models under selected dynamic symmetries.
Given a training dataset $\{(\textbf{x}^{(k)},y^{(k)})\}$ containing $N$ samples, the model is trained by minimizing the loss function